# Foundations of Cyber Resilience: The Confluence of Game, Control, and Learning Theories

### v1

@@ 1 Introduction to Cyber Resilience
Cyber resilience of a network or organization refers to its ability to prepare for, respond to, and recover from cyber threats or incidents. It represents a departure from classical cybersecurity measures, including cryptography, firewalls, and intrusion detection systems, which aim to prevent attackers from infiltrating or causing harm to the network. Instead, cyber resilience focuses on the capacity to respond to and recover from threats. Cyber resilience and cybersecurity are complementary to each other. Despite efforts and investments in cybersecurity for protection, perfect security cannot be guaranteed. There will always be threats that cannot be fully mitigated. These threats may include zero-day attacks, which are unknown to the network, or black swan events that are costly to secure against. Therefore, with limited resources and budgets, networks must decide how to protect themselves from threats to ensure the continuity of their applications and missions.
Resilience provides a complementary approach, as illustrated in Figure 1. When threats are difficult to prevent successfully, there is a need to find ways to respond to them after they have gained initial access to the network, deter further infiltration by attackers, and recover from any damage incurred. Additionally, it is important to recognize that the success of security mechanisms is probabilistic. A higher probability of success will require a greater allocation of resources. Despite the careful deployment of cybersecurity measures, there will still be a probability, albeit small, for attackers to succeed in bypassing cryptographic protocols and firewalls to gain access to the network.
@@
Within the set of exploitable vulnerabilities, we categorize them into two types of risks: inelastic risk and elastic risk. Elastic risks are those triggered by events that can only be mitigated elastically through resilience strategies. Conversely, inelastic risks are triggered by events that can be prevented with high probability using security mechanisms.
Figure 2: The vulnerabilities and their associated threats on the attack surface are represented in two dimensions: their exploitability and their impact. They span across cyber, physical, and human layers of networks or organizations. These vulnerabilities are categorized into two types of risks. One type is inelastic risks, which are managed by defensive security mechanisms, while the other type is elastic risks, which are addressed through resilience mechanisms.
@@ 3 Network and Dynamic Effects on Cyber Risks
A large network can consist of multiple interconnected subnetworks. The risk within one subnetwork can be influenced by others, as they are interconnected. A threat to one subnetwork can lead to consequences within that subnetwork and may also impact other connected subnetworks. As illustrated in Figure [influence], the states of other subnetworks are aggregated as $\Sigma_{t}$, and the well-being of these other subnetworks can affect the state of the subnetwork $X_{t}$. The interdependencies among subnetworks pose a significant challenge in managing cyber risks. Risk in one subnetwork can propagate to other subnetworks and vice versa. Such interdependencies have been vastly investigated in critical infrastructure networks [9, 29, 80, 28, 81] and communication networks [16, 15, 18, 34]. Defending against threats originating from within the network itself is challenging, and preparing for risks from connected networks is even more complex. Subnetworks may face threats from other networks for which they are not directly prepared. Therefore, resilience is critical in addressing these challenges, particularly due to the unknown risks that propagate through network connections.
@@ 4 Cyber Resilience and Cyber Risk Metrics
Numerous metrics can be proposed to quantify resilience. For instance, the total losses incurred between $t_{2}$ and $t_{5}$, denoted by $J_{E}$, reflect the impact of the event initiated by the attacker. Other metrics include $M$ and $D$, which gauge the severity of damage and the effectiveness of recovery in the responsive stage. Timing is also crucial; response time $t_{4}-t_{2}$ indicates how quickly recovery measures take effect, while retrospective time $t_{5}-t_{2}$ reflects the time taken for the system to rectify the problem retrospectively.
@@ 5.2 Game-theoretic foundations
Equilibrium analysis of dynamic games within this forward-looking context often employs dynamic programming principles or moving-horizon schemes, which are fundamental control-theoretic techniques utilized to tackle long-term problems. The control framework introduced in 1 is extended to a dynamic game framework with a more intricate modeling of adversarial behaviors, thereby creating a multi-player control problem. Such analysis yields a risk trajectory of the network environment in the face of adversaries. This analysis facilitates the measurement of network risks based on the prescribed attack model. The sequence of stage-by-stage risks corresponds to the cyber resilience metric depicted in Figure XX. In essence, at each stage, a game can assess risks by predicting the consequences of interactions between threats and defenses.
@@ 6 Future Directions
Incorporating these emerging methods into cybersecurity research and practice holds great promise for advancing our understanding of cyber threats and improving our ability to defend against them. By leveraging nonequilibrium learning, large language models, and foundation models with strategic learning, organizations can stay ahead of evolving threats and improve their cyber resilience in an increasingly complex and dynamic threat landscape.
@@ 7 Conclusions
The confluence of game, control, and learning theories emerges as a cornerstone for cyber resilience, offering a multifaceted approach to combat the ever-evolving landscape of cyber threats. By integrating insights from game theory to model strategic interactions, control theory to design dynamic defense mechanisms, and learning theory to adapt and evolve in response to uncertainties, organizations can improve their cyber resilience posture. The proactive, responsive, and retrospective stages of cyber resilience underscore the need for a holistic and adaptive approach to cybersecurity, where continuous learning and strategic decision-making play pivotal roles in safeguarding networks and critical assets. As cyber threats continue to escalate in complexity and frequency, the interdisciplinary framework presented in this chapter serves as a roadmap for organizations seeking to enhance their cyber resilience capabilities and navigate the intricate dynamics of modern cybersecurity challenges.

### v2

@@ 1 Introduction to Cyber Resilience
Cyber resilience of a network or organization refers to its ability to prepare for, respond to, and recover from cyber threats or incidents [24, 92, 72, 73]. It represents a departure from classical cybersecurity measures, including cryptography, firewalls, and intrusion detection systems, which aim to prevent attackers from infiltrating or causing harm to the network. Instead, cyber resilience focuses on the capacity to respond to and recover from threats. Cyber resilience and cybersecurity are complementary to each other. Despite efforts and investments in cybersecurity for protection, perfect security cannot be guaranteed. There will always be threats that cannot be fully mitigated. These threats may include zero-day attacks, which are unknown to the network, or black swan events that are costly to secure against. Therefore, with limited resources and budgets, networks must decide how to protect themselves from threats to ensure the continuity of their applications and missions.
As illustrated in Figure 1, resilience provides a complementary approach [97, 105, 48]. When threats are difficult to prevent successfully, there is a need to find ways to respond to them after they have gained initial access to the network, deter further infiltration by attackers, and recover from any damage incurred. Additionally, it is important to recognize that the success of security mechanisms is probabilistic. A higher probability of success will require a greater allocation of resources. Despite the careful deployment of cybersecurity measures, there will still be a probability, albeit small, for attackers to succeed in bypassing cryptographic protocols and firewalls to gain access to the network.
@@
Within the set of exploitable vulnerabilities, we categorize them into two types of risks: avertable risk and elastic risk. Elastic risks are those triggered by events that can only be mitigated elastically through resilience strategies. Conversely, avertable risks are triggered by events that can be prevented with high probability using security mechanisms.
Figure 2: The vulnerabilities and their associated threats on the attack surface are represented in two dimensions: their exploitability and their impact. They span across cyber, physical, and human layers of networks or organizations. These vulnerabilities are categorized into two types of risks. One type is avertable risks, which are managed by defensive security mechanisms, while the other type is elastic risks, which are addressed through resilience mechanisms.
@@ 3 Network and Dynamic Effects on Cyber Risks
A large network can consist of multiple interconnected subnetworks. The risk within one subnetwork can be influenced by others, as they are interconnected. A threat to one subnetwork can lead to consequences within that subnetwork and may also impact other connected subnetworks. As illustrated in Figure 4, the states of other subnetworks are aggregated as $\Sigma_{t}$, and the well-being of these other subnetworks can affect the state of the subnetwork $X_{t}$. The interdependencies among subnetworks pose a significant challenge in managing cyber risks. Risk in one subnetwork can propagate to other subnetworks and vice versa. These interdependencies can arise from three sources. Firstly, threats can exploit the connectivity between subnetworks, making it more convenient for attackers to deploy threats. For instance, attackers may compromise neighboring subnetworks before targeting the intended one. In this way, the threat space $T_{t}$ is expanded for the targeted network as threats to connected networks also affect its security. Secondly, vulnerabilities in the connections with neighboring subnetworks bring in new vulnerabilities of the subnetwork itself. Additionally, vulnerabilities in the nodes connected to it become potential threats that the subnetwork must address. In this way, the vulnerability space $V_{t}$ is expanded for the targeted network. Thirdly, interdependencies can arise through consequences. Risks can propagate through state influence, where the state of one subsystem influences others through physical dynamics. In this way, the consequence $X_{t}$ or $Y_{t}$ is directly affected.
Such interdependencies have been vastly investigated in critical infrastructure networks [12, 35, 106, 34, 107] and communication networks [21, 20, 23, 41]. Defending against threats originating from within the network itself is challenging, and preparing for risks from connected networks is even more complex. Subnetworks may face threats from other networks for which they are not directly prepared. Therefore, resilience is critical in addressing these challenges, particularly due to the unknown risks that propagate through network connections.
@@ 4 Cyber Resilience and Cyber Risk Metrics
Numerous metrics can be proposed to quantify resilience, which can be classified into three categories. The first category is performance-based metrics. These metrics require the monitoring of system states and provide accurate measures of resilience when credible information about system states is available. For example, the total losses incurred between $t_{2}$ and $t_{5}$, denoted by $J_{E}$, reflect the impact of the event initiated by the attacker. Other metrics in this category include $M$ and $D$, which assess the severity of damage and the effectiveness of recovery during the responsive stage. The second category comprises latency-based metrics, which are dependent on timing. These metrics measure the speed or frequency of reaction against risks. For instance, the response time $t_{4}-t_{2}$ indicates how quickly recovery measures take effect, while the retrospective time $t_{5}-t_{2}$ reflects the time taken for the system to rectify the problem retrospectively. The third category consists of knowledge-based metrics, which are associated with the known area of the attack surface (AS). Resilience in this context depends on the comprehensiveness of the knowledge about the system. For example, the percentage of the system model that can be expressed deterministically.
@@ 5.2 Game-theoretic foundations
Equilibrium analysis of dynamic games within this forward-looking context often employs dynamic programming principles or moving-horizon schemes, which are fundamental control-theoretic techniques utilized to tackle long-term problems. The control framework introduced in 1 is extended to a dynamic game framework with a more intricate modeling of adversarial behaviors, thereby creating a multi-player control problem. Such analysis yields a risk trajectory of the network environment in the face of adversaries. This analysis facilitates the measurement of network risks based on the prescribed attack model. The sequence of stage-by-stage risks corresponds to the cyber resilience metric depicted in Figure 8. In essence, at each stage, a game can assess risks by predicting the consequences of interactions between threats and defenses.
@@ 6 Future Directions
The consequences of cyber risks in critical infrastructures and industrial networks often extend to risks within physical systems. Consequently, the risk is not solely confined to the cyber domain but incorporates both cyber and physical domains, making it a cyber-physical challenge. Thus, associated resilience mechanisms must not only address cyber threats but also consider the implications for physical systems. Therefore, the design of resilience mechanisms needs to be physics-informed, i.e., integrating insights from physics to ensure compliance with operational level constraints. Additionally, it is essential to develop physical-level resilient mechanisms, such as controllers, to mitigate impacts on physical systems. Consequently, cyber resilience must be jointly designed with physical resilience, emphasizing a co-design paradigm for networks [105, 39, 80, 82, 36]. This approach is crucial and has been applied in understanding cyber-physical resilience across various domains, including nuclear systems [91, 90, 75], power systems [96, 97, 10, 94, 104], industrial control systems [15, 14, 89, 81, 79, 20], multi-agent systems [92, 56, 72, 11, 56, 57], and interdependent infrastructures [12, 9, 35, 36, 74].
Incorporating emerging methods into cybersecurity research and practice holds great promise for advancing our understanding of cyber threats and improving our ability to defend against them. By leveraging nonequilibrium learning, large language models, and foundation models with strategic learning, organizations can stay ahead of evolving threats and improve their cyber resilience in an increasingly complex and dynamic threat landscape.
@@ 7 Conclusions
Dealing with cyber risk in today’s world is increasingly challenging. Mitigating cyber risk should not only involve cyber security defense mechanisms but also cyber resilience mechanisms, which complement each other to strengthen the cyber risk posture of the network. Understanding cyber risk and the associated resilience mechanisms requires the development of fundamental principles, metrics, and tools to analyze and design networks to achieve resilience goals. This chapter introduces game theory, control theory, and learning theory as the foundation for cyber resilience. Game theory has been shown to be suitable for modeling strategic interactions. Control theory has been pivotal in dealing with dynamics and designing dynamic defenses. Learning theory plays an important role in bridging theory and practice through autonomous adaptation and dealing with uncertainties. The confluence of game theory, control theory, and learning theory emerges as a cornerstone for cyber resilience, enabling organizations to improve their cyber resilience posture. As cyber threats continue to escalate in complexity and frequency, the interdisciplinary framework presented in this chapter serves as a roadmap for organizations seeking to enhance their cyber resilience capabilities and navigate the intricate dynamics of modern cybersecurity challenges.